OpenStudy (anonymous):

-4x-15y=-17 -x+5y=-13 i have to find x and y by elimination please show work !!

jimthompson5910 (jim_thompson5910):

First multiply both sides of equation 2 by 3 to get -x+5y=-13 3(-x+5y) = 3(-13) -3x + 15y = -39 --------------------------- So we now have this new system -4x - 15y = -17 -3x + 15y = -39 Now add the two equations to get this -4x - 15y = -17 -3x + 15y = -39 --------------- -7x + 0y = -56 which means -7x = -56 ----> x = ???

OpenStudy (anonymous):

8??

jimthompson5910 (jim_thompson5910):

good x = 8 now plug that into -x+5y=-13 and solve for y

OpenStudy (anonymous):

-1??

jimthompson5910 (jim_thompson5910):

perfect so x = 8 and y = -1

OpenStudy (anonymous):

awesoommmmmeee is it possible that i cud multiply both equations by a number

jimthompson5910 (jim_thompson5910):

you could, but it would take more work also, the numbers would have to be different

OpenStudy (anonymous):

oo i c